# AI for Just Work: Constructing Diverse Imaginations of AI beyond “Replacing Humans”

### v1

@@ 4.3 Applying the imagination: a case study
The foundational assumptions and AI properties can be applied and extended in different contexts and subfields of AI. Next, we showcase an example that demonstrates the process of applying the new imagination of AI in research and potential applications of the medical image synthesis (MISyn) task. The MISyn task is to generate “visually realistic and quantitatively accurate images” in biomedicine (Frangi et al., 2018). Our recent work “Ethical Medical Image Synthesis” (Jin et al., 2025) conducts a thorough ethical analysis and proposes ethical criteria for MISyn that align with the new imagination of “AI for just work.” We highlight key points that link the imagination to corresponding actions in MISyn research and peer review.
Based on A1, we identify the intrinsic limits of MISyn, including that synthetic images are not automatically grounded in real medical phenomena unlike medical images. Ignoring this limit permits the misinformation of MISyn that could cause harms to clinical workers and patients. To prevent the misinformation of MISyn, we propose several ethical criteria, including setting up technical standards, terminology, and usage declaration to clearly differentiate synthetic images from medical images in all outputs of MISyn techniques.
Based on P1 and A3, we propose the ethical criterion that the problem formulation of MISyn should ground in real-world medical problems. Otherwise, as revealed by our ethical analysis, it is easy to mistake the technical progress in the model space as the real progress in clinical problems, which only benefits technical sectors rather than non-technical stakeholders. We also propose the five-phased evaluation paradigm to ground the evaluation of MISyn in clinical settings.
The ethical criteria also emphasize the importance of conducting limitation analyses, and extend the three aspects of limitation assessment in P2 to tailor to the risks and limits of MISyn techniques. The right to reject AI (P3) and the protection of stakeholders’ rights (A2, P4) are embedded in our proposed non-technical checklist for stakeholders to question the appropriateness of MISyn. To facilitate the implementation of the ethical criteria in MISyn research and applications, we conduct paper reviews to critically analyze two MISyn works published in high-impact journals using the ethical criteria, and show concrete examples of the gap between existing practice and the ethical criteria guided by different imaginations. The detailed changes of practice are listed in Appendix A.
@@ 5 Conclusion
In this work, we explore the “why” question in AI. We point out flaws in the current mainstream imagination of AI, and demonstrate one possible way to reimagine AI for the flourishing of humanity and the planet. This work calls the community to critically examine and diversify our mindsets on the foundational assumptions of AI that are usually taken for granted. We hope this work can inspire more works to diversify our collective imaginations of AI for ethical purposes, and encourage collective efforts within and outside the community to create space and support for these diverse imaginations to grow and thrive.
Acknowledgements
We thank Prof. Wendy Hui Kyong Chun, Prof. Martin Ester, Adrian D’Alessandro, Ben Cardoen, Mark Nielsen, and Kumar Abhishek for their helpful comments and discussions.
@@ Acknowledgements
This work originates from our concern about the increasingly negative impacts of AI to society and the environment. In this work, we try to provide our diagnosis and treatment of the root causes of negative social impacts of AI by tackling the imagination of AI. While we call for positive changes in our imaginations of AI and hope this work can create positive impacts to the AI community, we are also aware that good intentions or imaginations alone do not automatically guarantee good impacts, a prominent example is the “ethics washing” phenomenon in AI (Rességuier & Rodrigues, 2020; Wagner, 2019; Ochigame, 2022). Therefore, we embed ethical values in the methodology and technical practice of AI, and combine it with power check mechanisms that strengthen the social and state power to check technical power, such as conducting critical reflections of our practice to reflect our responsibilities as researchers, scientists, and citizens, opening dialogues with the public, protecting workers’ rights to decide, and regulations. We hope implementation of these mechanisms can equip ethics with its teeth and steer towards more positive impacts of the diverse and ethical imaginations of AI to society.
@@ Appendix A Applying the imagination of “AI for just work”: Case study of the medical image synthesis task
In Section 4.3, we showcase our application of the imagination of “AI for just work” to the specific medical image synthesis (MISyn) task based on our recent work “Ethical Medical Image Synthesis” (Jin et al., 2025). We use this case because this is our first and currently the only project that applies the new imagination in research practice. To supplement this case, we detail the changes of practice before and after applying the imagination in LABEL:tab:case. In the “Ethical Medical Image Synthesis” paper, to provide a hands-on guidance on how to apply the proposed ethical criteria in research, we conducted paper reviews of two MISyn works published in high-impact journals in the medical image analysis field. The examples of practice in LABEL:tab:case are from the paper reviews.
The case study on the medical image synthesis task also indicates that even if the task is intended for good purposes, such as this task of AI for medicine, and even if the target domain of application has already established strong ethics in its practice, such as this target domain of medicine, the penetration of the mainstream imagination of AI (MIA) into technical practice can still pose unscientific and unethical risks (Zheng et al., 2024).
@@
2.Proposed power to reject AI We develop a checklist for non-technical stakeholders, regulators, and the public to check if a MISyn fulfills the ethical criteria, and enable them to question the appropriateness of the MISyn. It assumes that workers have the right to refuse AI (P3). It reflects the new imagination that the design of AI should try to enhance worker power (A2, P4).
@@
3.Proposed problem formulation We provide ethical criteria to ground MISyn techniques in clinical needs, and provide specific suggestions for the two MISyn works, such as grounding the need for synthetic images in clinical problems of wound segmentation when real data is extremely limited, and grounding the text-to-image generative feature in surgical simulation with user-customized text inputs. It assumes that the adaptation of an AI technique to a clinical problem does not happen automatically and requires considerable human efforts. Therefore, it is the technical sector’s responsibility to incorporate task-specific requirements into the technique design from the beginning. It reflects the new imagination to ground the technical problem formulation in real-world tasks (P1).
@@
4.Proposed evaluation reporting standard We provide the ethical criteria for technical and non-technical reviewers to check if the evaluation contains unscientific claims or overclaims. It assumes that the AI evaluation and reporting should maintain high standards of scientific validity and rigor as in other research fields. It reflects the new imagination that prioritizes ethical and scientific values over technical progress (A3).
@@
6.Proposed limitation analysis We propose the ethical criteria of three aspects of limitation analysis, including the thorough limitation assessment quantitatively and/or qualitatively on the scope, weaknesses, and failure modes of the proposed MISyn technique, acknowledging and declaring the limitations in technical design and assumptions, and acknowledging the intrinsic limits of MISyn techniques that cannot be overcome by technical advances. It assumes that AI has limits and weaknesses, and the role of thorough limitation analysis is equivalent to or even more important than the benefit analysis of AI. It reflects the new imagination that AI technologies have intrinsic limits and fundamental dependencies (A1).
Appendix B Bibliographic essay on the multidisciplinary knowledge background
@@
B.1 Section 1. The imagination of AI: What is AI for?
B.1.1 The concept of imagination in social science
Imagination, or social imagery, is a concept usually used in social science to describe “the set of values, institutions, laws, and symbols through which people imagine their social whole. It is common to the members of a particular social group and the corresponding society.” [^6] https://en.wikipedia.org/wiki/Imaginary_(sociology) It is “the web of meanings that binds a particular society together” (Earle, 2020) and provides “structures that guide our collective and individual actions and values.” [^7] https://socialimaginaries.org/the-imaginary-system-of-society/
In her book Imagination: A Manifesto (Benjamin, 2024), sociologist Ruha Benjamin outlines the role of collective imagination in social change, and she provides the definition of imagination in the Preface of the book as follows:

### v2

@@ 4.3 Applying the imagination: a case study
The foundational assumptions and AI properties can be applied and extended to different contexts, tasks, and subfields of AI. Next, we showcase an example that demonstrates the process of applying the new imagination of AI in research and potential applications of the medical image synthesis (MISyn) task. The MISyn task is to generate “visually realistic and quantitatively accurate images” in biomedicine [46]. Our recent work “Ethical Medical Image Synthesis” [62] conducts a thorough ethical analysis and proposes ethical practice recommendations for MISyn that align with the new imagination of “AI for just work.” We highlight key points that link the imagination to corresponding actions in MISyn research and peer review.
Based on A1, we identify the intrinsic limits of MISyn, including that synthetic images are not automatically grounded in real medical phenomena unlike medical images. Ignoring this limit permits the misinformation of MISyn that could cause harms to clinical workers and patients. To prevent the misinformation of MISyn, we propose several ethical practice recommendations, including setting up technical standards, terminology, and usage declaration to clearly differentiate synthetic images from medical images in all outputs of MISyn techniques.
Based on P1 and A3, we propose the ethical practice recommendation that the problem formulation of MISyn should ground in real-world medical problems. Otherwise, as revealed by our ethical analysis, it is easy to mistake the technical progress in the model space as the real progress in clinical problems, which only benefits technical sectors rather than non-technical stakeholders. We also propose the five-phased evaluation paradigm to ground the evaluation of MISyn in clinical settings.
The ethical practice recommendations also emphasize the importance of conducting limitation analyses, and extend the three aspects of limitation assessment in P2 to tailor to the risks and limits of MISyn techniques. The right to reject AI (P3) and the protection of stakeholders’ rights (A2, P4) are embedded in our proposed oversight recommendations for stakeholders to question the appropriateness of MISyn. To facilitate the implementation of the ethical practice recommendations in MISyn research and applications, we conduct paper reviews to critically analyze two MISyn works published in high-impact journals using the ethical practice recommendations, and show concrete examples of the gap between existing practice and the recommendations guided by different imaginations. Detailed changes in technical and non-technical practice are listed in Appendix A.
@@ 5 Conclusion
In this work, we explore the “why” question in AI. We point out flaws in the current social structures of the AI community, embodied in the form of the mainstream imagination of AI, and demonstrate one possible way to reimagine AI for the flourishing of humanity and the planet. In addition to this new imagination of “AI for just work,” there are various other possibilities and initiatives to construct more ethical imaginations of AI. Examples include the Indigenous AI and Abundant Intelligences programs [70, 69], the First Languages AI Reality initiative [78], the Common Voice project [80], among many others, which usually appear in subcommunities such as AI for (Social) Good, AI for Science, FAccT, and AIES (AI, Ethics, and Society). Our work calls the AI community to critically examine and diversify our mindsets on the foundational assumptions of AI that are usually taken for granted, and expand the influence of more ethical imaginations from subcommunities to the mainstream AI development agendas and techniques. To monitor and track structural changes in improving the degree of ethics and diversity of the community’s collective imaginations, future work can conduct empirical studies and explore methodologies of assessment, and we provide a preliminary discussion in Appendix B for reference. We hope this work can inspire more works to diversify our collective imaginations of AI for more ethical purposes, and encourage collective efforts within and outside the community to create space and support for these diverse imaginations to grow, thrive, and become a central component of the AI community.
Acknowledgments and Disclosure of Funding
We thank Prof. Wendy Hui Kyong Chun, Prof. Martin Ester, Adrian D’Alessandro, Dr. Ben Cardoen, Mark Nielsen, and Kumar Abhishek for their helpful comments and discussions.
@@
This work originates from our concern about the increasingly negative impacts of AI to society and the environment. In this work, we try to provide our diagnosis and treatment of the root causes of negative social impacts of AI by tackling the imagination of AI. While we call for positive changes in our imaginations of AI and hope this work can create positive impacts to the AI community, we are also aware that good intentions or imaginations alone do not automatically guarantee good impacts, a prominent example is the “ethics washing” phenomenon in AI [94, 103, 85]. Therefore, we embed ethical values in the methodology and technical practice of AI, and combine it with power check mechanisms that strengthen the social and state power to check technical power, such as conducting critical reflections of our practice to reflect our responsibilities as researchers, scientists, and citizens, opening dialogues with the public, protecting workers’ rights to decide, and regulations. We hope implementation of these mechanisms can equip ethics with its teeth and steer towards more positive impacts of the diverse and ethical imaginations of AI to society.
@@ Appendix A Applying the imagination of “AI for just work”: Case study of the medical image synthesis task
In Section 4.3, we showcase our application of the imagination of “AI for just work” to the specific medical image synthesis (MISyn) task based on our recent work “Ethical Medical Image Synthesis” [62]. As we emphasized in the paper, imaginations and values shape how we frame, design, and evaluation AI techniques. Therefore, our actions for change in our collective imaginations can influence extensive technical practice of different tasks at hand. We use this case because this is our first and currently the only project that applies the new imagination in research practice. To supplement this case, we detail the changes of practice before and after applying the imagination in LABEL:tab:case. In the “Ethical Medical Image Synthesis” paper, to provide a hands-on guidance on how to apply the proposed ethical practice recommendations in research, we conducted paper reviews of two MISyn works published in high-impact journals in the medical image analysis field. The examples of practice in LABEL:tab:case are from the paper reviews.
The case study on the medical image synthesis task also indicates that even if the task is intended for good purposes, such as this task of AI for medicine, and even if the target domain of application has already established strong ethics in its practice, such as this target domain of medicine, the penetration of the mainstream imagination of AI (MIA) into technical practice can still pose unscientific and unethical risks [112].
@@
2.Proposed power to reject AI We develop a checklist for non-technical stakeholders, regulators, and the public to check if a MISyn fulfills the ethical practice recommendations, and enable them to question the appropriateness of the MISyn. It assumes that workers have the right to refuse AI (P3). It reflects the new imagination that the design of AI should try to enhance worker power (A2, P4).
@@
3.Proposed problem formulation We provide ethical practice recommendations to ground MISyn techniques in clinical needs, and provide specific suggestions for the two MISyn works, such as grounding the need for synthetic images in clinical problems of wound segmentation when real data is extremely limited, and grounding the text-to-image generative feature in surgical simulation with user-customized text inputs. It assumes that the adaptation of an AI technique to a clinical problem does not happen automatically and requires considerable human efforts. Therefore, it is the technical sector’s responsibility to incorporate task-specific requirements into the technique design from the beginning. It reflects the new imagination to ground the technical problem formulation in real-world tasks (P1).
@@
4.Proposed evaluation reporting standard We provide the ethical practice recommendations for technical and non-technical reviewers to check if the evaluation contains unscientific claims or overclaims. It assumes that the AI evaluation and reporting should maintain high standards of scientific validity and rigor as in other research fields. It reflects the new imagination that prioritizes ethical and scientific values over technical progress (A3).
@@
6.Proposed limitation analysis We propose the ethical practice recommendations of three aspects of limitation analysis, including the thorough limitation assessment quantitatively and/or qualitatively on the scope, weaknesses, and failure modes of the proposed MISyn technique, acknowledging and declaring the limitations in technical design and assumptions, and acknowledging the intrinsic limits of MISyn techniques that cannot be overcome by technical advances. It assumes that AI has limits and weaknesses, and the role of thorough limitation analysis is equivalent to or even more important than the benefit analysis of AI. It reflects the new imagination that AI technologies have intrinsic limits and fundamental dependencies (A1).
Appendix B Methods of evaluating the ethics and diversity of collective imaginations
With the implementation of structural changes to improve the ethics and diversity of collective imaginations, future work can conduct empirical studies to monitor and track the effects of implementation. We provide a preliminary discussion on the methodologies of assessing the ethics and diversity of collective imaginations in AI projects/works/papers, which may involve the following three steps.
  1. M1
    Identify a collection of samples.
    Like the literature review method, we first need to identify the target samples of AI works. Depending on the specific research questions, the samples can be a collection of papers from certain venues, as Birhane et al., 2022b [22] did, or the most cited papers to reflect a certain group tendency, as Birhane et al., 2022a [20] did.
  2. M2
    For each sample, identify its embedded values or imaginations, assess whether it is ethical, if so, identify its ethical direction.
    The embedded values or imaginations can usually be identified from its motivation, problem formulation, main claims, contribution statements, research questions, hypothesis, or evaluation objectives.
    To judge whether the work embedded ethical imaginations, we can define several ethical assessment features to annotate paper. A paper that shows one of several features to a certain degree can be regarded as fulfilling the ethical criteria to have ethical imaginations.
    The definition of ethical assessment features should encompass diverse directions of ethical imaginations. The ethical assessment features may not be purely based on the paper’s claim or intended goal (for example, AI for healthcare/science/social good, fair/responsible/explainable AI, or human-centered/human-augmenting AI), but also need to consider whether the ethical imaginations have been really incorporated in the paper’s detailed practice. Otherwise, the claimed ethical goal can easily become “ethics washing” for the actual non-ethical or unethical practice in technical design and evaluation details. Our new imagination of “AI for just work” depicts some features of the foundational assumptions A1-A3 and AI properties P1-P4.
    For reference, Birhane et al., 2022a [20] and Birhane et al., 2022b [22] provide the full methodological pipelines to define and annotate papers based on AI ethical features or values encoded in AI projects.
    For works that fulfill the ethical criteria, their ethical direction should be annotated to document what kinds of welfare the work aims to address.
  3. M3
    Assess diversity of the collection of samples regarding the ethical direction.
    From the annotated ethical directions of the collection of samples, we can cluster them and use some quantitative metrics to measure the degree of diversity regarding the ethical directions.
We note that ethics is a complex concept, and any working definitions and assessment features of ethics cannot be complete and depend on particular perspectives that frame the assessment. Alexandrova and Fabian, [11] suggest a bottom-up participatory research process to incorporate diverse perspectives. The method and procedure of assessment should be transparent and well documented, so that the value judgments in the assessment can fairly answer the research questions and withstand scrutiny.
Appendix C Glossary
To enable readers to better refer to and understand the interdisciplinary terminologies introduced in our discussion, we provide a Glossary in LABEL:tab:glossary that lists the relevant terminologies, their definition, relevant technical examples to instantiate them, their locations in the paper, and references.
Table 2: Glossary of the terms introduced in the paper (in alphabetical order)
Term Definition Example Location in the paper Reference
Blind spot worldview It is a particular worldview that “had been grafted onto science.” Frank et al., [47] name it to denote “the dominant philosophical conception of science led to elevating mathematical abstractions as what is truly real and to devaluing the world of immediate experience,” causing the failure to see direct experience that makes science possible. We denote its opposite as the normal worldview in the paper. The manifestation of the blind spot worldview in AI is named by Frank et al., [47] as the “computational blind spot” that prevents us “from seeing how our experience of the world relates to our computational constructions.” S1 in Section 2.2, Fig. 2 Frank et al., [47]. More information in Section D.2.1
Co-production It depicts the relationship between technology and society, that technology acting as a mediator: technology is shaped by social factors, and technology in turn shapes the social arrangement. Racial and gender biases can be baked into the training data and loss functions of the AI model, and the deployed model influences how gender and race are perceived in society. S4 in Section 2.2 Matthewman, [74], Boenig-Liptsin et al., [24]
Complex system Complex systems are “Co-evolving multilayer networks” that are context-dependent and are composed of many non-linearly interacting elements. Examples of complex systems include language, human behavior, human mind, a living cell, financial market, social or natural phenomena. A1 in Section 4.1 Thurner et al., [102]
Displacement effect It is when capital takes over tasks previously performed by labor. Industrial robots replace human workers in assembly lines. S3 in Section 2.2 Acemoglu and Restrepo, 2019b [5]
Epistemic injustice It is a kind of injustice that “wrongs someone in their capacity as a subject of knowledge, and thus in a capacity essential to human value.” Claiming that AI can replace humans is a manifestation of epistemic injustice that downplays the epistemic footing of human experience in building and using AI systems. S2 in Section 2.2 Fricker, [50], Symons and Alvarado, [101]
Imagination Imagination, or social imagery, is a concept usually used in social science to describe “the set of values, institutions, laws, and symbols through which people imagine their social whole. It is “the web of meanings that binds a particular society together” [42] and provides “structures that guide our collective and individual actions and values.” The imagination of AI provides sense of meanings to answer the “why” question in AI, such as “why do we develop AI? What purposes does AI serve?” It is the general vision, frame, and mindset we have for the prospects of AI technologies. Section 1 Benjamin, [18], Earle, [42]. More information in Section D.1.1
Mainstream imagination of AI, MIA It denotes the dominant collective imaginations of the AI community. The current MIA emphasizes progress, performance, replacing humans and improving productivity. Section 1, Section 2.1 This is a term we introduce in the paper
Marginal productivity It is “the additional contribution that one more worker brings by increasing production or by serving more customers.” Acemoglu and Johnson, [2] provide an example that “when a car company introduces a better vehicle model, … this tends to increase the demand for the company’s cars, and both revenues per worker and worker marginal productivity rise. After all, the company needs more workers, such as welders and painters, to meet the additional demand, and it will pay them more, if necessary.” S3 in Section 2.2 Acemoglu and Johnson, [2]
Medical image synthesis A task in the medical image analysis domain that is to generate “visually realistic and quantitatively accurate images” in biomedicine An example of the task is to generate images to augment data when training medical imaging models. Section 4.3 Frangi et al., [46]
Paths of least resistance It is the metaphorical pathway that provides the least resistance for people to participate in social situations, among a set of alternative possibilities. The trendy research topics and mainstream research paradigms provide AI researchers the paths of least resistance in framing a research project, among many other potential research approaches, because trendy topics are more likely to gain attention and influence in the community, and mainstream paradigm is less likely to be questioned or challenged. Section 2.1, Section 3 Johnson, [63], Wikipedia
Power Power is “the asymmetric capacity to structure or alter the behavior of others.” People have the power to reject AI decisions is an example of power to alter the default behavior of the AI agent. S2 and S3 in Section 2.2 Boenig-Liptsin et al., [24]
Power structure It is “the distribution of power in a social system.” Currently, the technical stakeholders have more control in shape AI research agenda compared to the public, i.e., the technical stakeholders are dominant over the public in the power structure. When mechanisms have been implemented to empower the public to control over how AI can be developed, we say the power structure is shifted towards the public side. Section 3 Johnson, [63]
Productivity, or average productivity It is a worker’s average output. Acemoglu and Johnson, [2] provide an example that when industrial robots are installed, the company’s average productivity increases significantly. S3 in Section 2.2 Acemoglu and Johnson, [2]
So-so automation, or excessive automation It is when the productivity gain from automation is not enough to offset the displacement effect. Acemoglu and Johnson, [2] provide an example of the self-checkout kiosks in grocery stores: they “bring limited productivity benefits because they shift the work of scanning items from employees to customers. When self-checkout kiosks are introduced, fewer cashiers are employed, but there is no major productivity boost to stimulate the creation of new jobs elsewhere. Groceries do not become much cheaper, there is no expansion in food production, and shoppers do not live differently.” S3 in Section 2.2 Acemoglu and Restrepo, 2019c [6], Acemoglu and Restrepo, 2019a [4]
Social structure It is “the patterns of relationships and distributions that characterize the organization of a social system. Relationships connect various elements of a system (such as social statuses) to one another and to the system itself. Distributions include valued resources and rewards, such as power and income, and the distribution of people among social statuses. Structure can also refer to relationships and distributions among systems.” The social relationship of reviewer and author is a social structure that depicts people’s likely behavior when participating in the social system of peer review. Section 2.1 Johnson, [63]
Social system It is “an interconnected collection of social structural relationships, ecological arrangements, cultural symbols, ideas, objects, and population dynamics and conditions that combine to form a whole. Complex systems comprise smaller systems that are related to one another and the larger system through cultural, structural, ecological, and population arrangements and dynamics.” In the paper, we use the words “community”, “AI community”, or “research community” to denote the meaning of social system. Section 2.1 Johnson, [63]
Technological determinism “Technological determinism is a reductionist theory in assuming that a society’s technology progresses by following its own internal logic of efficiency, while determining the development of the social structure and cultural values.” In Section 3, we argue against an assumption that the internal logic of AI determines its development direction. With this assumption, the harms and benefits of AI development are regarded as inevitable consequences. Section 3 Wikipedia
Technological fix, or techno-solutionism “A technological fix, technical fix, technological shortcut or (techno-)solutionism is an attempt to use engineering or technology to solve a problem (often created by earlier technological interventions).” Selbst et al., [99] defines the “Solutionism Trap‘” as the “failure to recognize the possibility that the best solution to a problem may not involve technology.” An example is seen in S1 in Section 2.2 that whenever there are problems such as social or scientific problems, or problems caused by technologies, we resort to AI technology to solve them. S4 in Section 2.2 Wikipedia, [108], Morozov, [79]
Work, labor, Worker Borrowed from critical theory, “work” in the new imagination of “AI for just work” is defined as “meaningful, fulfilling activities” as the “social means necessary to live a flourishing life.” Work under this assumption includes a wider range of human activities that may not count as “work” in the current narrowly framed economic system based on market exchange. We use the words “work” and “labor” interchangeably. In the context of AI, workers include a wide range of people who work for or behind AI (such as data labor), who work with AI (such as users), and whose work or life is impacted by AI (such as stakeholders). A2 in Section 4.1 Wright, [109, 110]
Appendix D Bibliographic essay on the multidisciplinary knowledge background
@@
D.1 Section 1. The imagination of AI: What is AI for?
D.1.1 The concept of imagination in social science
Imagination, or social imagery, is a concept usually used in social science to describe “the set of values, institutions, laws, and symbols through which people imagine their social whole. It is common to the members of a particular social group and the corresponding society.” [^8] https://en.wikipedia.org/wiki/Imaginary_(sociology) It is “the web of meanings that binds a particular society together” [42] and provides “structures that guide our collective and individual actions and values.” [^9] https://socialimaginaries.org/the-imaginary-system-of-society/
In her book Imagination: A Manifesto [18], sociologist Ruha Benjamin outlines the role of collective imagination in social change, and she provides the definition of imagination in the Preface of the book as follows: